OpenStudy (anonymous):

If you know how to complete a square please help! In the function f(x) = 5(x^2 - 4x __) + 15, what number belongs in the blank to complete the square?

OpenStudy (anonymous):

x2−4x+? (b/2) (-4/2)^2 = +4 x^2-4x+4 The answer would be 4.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

ur welcome medal plz?

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

@clanger555 can you prove it your answer please that is right ?

OpenStudy (anonymous):

sure

jhonyy9 (jhonyy9):

thank you in advance

OpenStudy (anonymous):

x2−4x+? (b/2) (-4/2)^2 = +4 x^2-4x+4 The answer would be 4.

jhonyy9 (jhonyy9):

step by step please - i dont understand it so

jhonyy9 (jhonyy9):

f(x) = 5(x^2 - 4x __) + 15 how you make from these a square perfect ?

OpenStudy (anonymous):

@clanger555 Could you help with another one?

jhonyy9 (jhonyy9):

just bc. this is the exercise ,the question

OpenStudy (anonymous):

Thank you so much!